# PSHuman: Photorealistic Single-view Human Reconstruction using Cross-Scale Diffusion

### v1

@@ 3.2 SMPLX-initialized explicit human carving
Following the generation of multi-view color and normal images, we elaborate on our proposed SMPLX-initialized human carving module (Fig. 6) to obtain the textured 3D mesh.
@@
Numerous methodologies have been developed to leverage normal cues for human reconstruction. However, a significant proportion of them employ implicit functions (e.g. MLP) to map the normal feature as implicit surfaces. This process, while effective in certain scenarios, often results in a lack of fine geometric details. Even with BiNI used in ECON, the overall geometry still exhibits a notable degradation. Taking advantage of the multi-view consistent normal maps, we opt to fuse it directly with the explicit triangle mesh. Our reconstruction module consists of three main stages: SMPL-X initialization, differentiable remeshing, and appearance fusion.
@@
SMPL-X initialization. The process commences with human mesh initialization, utilizing the aforementioned SMPL-X estimation, which provides a strong body prior, effectively mitigating unnecessary face pruning and densification during subsequent geometry optimization. However, it is noteworthy that the generated multiple views may exhibit slight misalignment with the SMPL model due to normalization and recentering procedures tailored for the diffusion model. Drawing inspiration from ICON, we optimize SMPL-X’s translation, shape, and pose by minimizing the pixie-aligned error of multi-view normal and silhouette. The alignment process is computationally efficient, typically requiring only seconds to complete.
@@
Remeshing with differentiable rasterization. Given the initial human prior, we utilize differentiable rasterization to carve the details based on observational normal maps. While a common approach involves adding per-vertex displacement to the coarse canonical mesh, this method encounters difficulties when modeling complex details, such as loose clothing. To address this limitation, we directly optimize the SMPL topology, encompassing both vertex positions $V$ and face edges $F$. The optimization procedure iteratively applies vertex displacement and remeshing to the triangle mesh, utilizing the optimizer proposed in (Palfinger, 2022). The optimization objective can be written as
$\tilde{V},\tilde{F}=\operatorname*{arg\,min}_{V,F}\sum_{i=1}^{N}w_{i}(\|N_{i}-\hat{N}_{i}\|_{2}+\|S_{i}-\hat{S}_{i}\|_{2})+\lambda\sum_{j}(n_{j}-n_{j}^{\text{neig}})$ (7)
@@
where $n_{j}$ and $n_{j}^{\text{neig}}$ denote the vertex normal and the average normal of neighboring vertices, The regularization weight $\lambda$ is set to $0.02$. We execute $700$ optimization steps to achieve optimal performance. Following the mesh optimization, we employ Poisson reconstruction Kazhdan & Hoppe (2013) to complete minor invisible areas, such as the chin. Additionally, we offer the option to substitute the hands with the estimated SMPL-X results (Xiu et al., 2023).
@@
Appearance fusion. Upon obtaining the 3D geometry, our objective is to derive the high-fidelity texture matching the reference image. Despite the availability of multi-view images, direct projection onto the mesh results in conspicuous artifacts, arising from the cross-view inconsistency and inaccurate foreground segmentation. To overcome this, we perform texture fusion and optimize the per-vertex color by minimizing the view-dependent MSE loss between the rendered color images and generated ones utilizing differentiable rendering. Finally, we compute a visibility mask and perform topology-aware interpolation to complete the minor unobserved area.
@@ 4 Experiments
Dataset. We conduct experiments on widely used 3D human datasets, including high-quality human scans (THuman2.1 Yu et al. (2021b) and CustomHumans Ho et al. (2023)) captured with a dense DSLR rig and temporal sequence of scans (CAPE Ma et al. (2020)) captured with a body scanner. Specifically, our training dataset comprises $2,385$ scans from THuman2.1 and $647$ scans from CustomHumans. These datasets are selected due to their provision of SMPL-X parameters. For quantitative evaluation, we utilize the remaining $60$ scans from THuman2.1 and $150$ scans from CAPE, with CAPE being subdivided into “CAPE-FP” and “CAPE-NFP” to assess generalization on real-world scenarios. Additionally, we curate a selection of cases from the Internet and SHHQ Fu et al. (2022) fashion data for qualitative comparison.
Metric. To assess reconstruction capability, we employ three primary metrics: 1-directional point-to-surface (P2S), $L_{1}$ Chamfer Distance (CD), and Normal Consistency (NC). CD and P2S quantify the distance between predicted and ground-truth meshes, while NC measures the cosine distance between surface normals. For appearance quality evaluation, we utilize peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and learned perceptual image patch similarity (LPIPS).
Table 1: Quantitative comparison of geometry quality. To avoid the impact of pose estimation errors on the evaluation, ground-truth SMPL-X models are used during testing. The units for Chamfer and P2S are in cm. The top two results are colored as first second.
@@
Method Publication Cham. Dist $\downarrow$ P2S $\downarrow$ NC $\uparrow$ Cham. Dist $\downarrow$ P2S $\downarrow$ NC $\uparrow$ Cham. Dist $\downarrow$ P2S $\downarrow$ NC $\uparrow$
@@ 4.1 Comparisons
Baselines. We conducted a comprehensive comparison of our method against state-of-the-art single-view human reconstruction approaches, including PIFu Saito et al. (2019), PIFuHD Saito et al. (2020), PaMIR Zheng et al. (2021), ICON Xiu et al. (2022), ECON Xiu et al. (2023), GTA Zhang et al. (2024a), SiFU Zhang et al. (2024b), and SiTH Ho et al. (2024). For SMPL-based methods, we utilize PIXIE Yu et al. (2021a) for estimation. We also report the results with ground-truth SMPL-X to isolate the impact of pose estimation errors.
Comparison of geometry quality. Our method demonstrates superior geometric quality compared to existing approaches, particularly without an SMPL-X body prior (Tab. 1). Unlike template-based methods, which are susceptible to SMPL-X prediction errors, our method supports template-free training, thereby offering enhanced generalization capability. When incorporating the body prior, our method consistently outperforms previous works, demonstrating unprecedented accuracy on complex posed humans. The qualitative comparison in Fig. 2 also showcases the superiority of PSHuman, featuring with complete shape, detailed face and natural-looking clothing folds.
@@
Comparison of appearance quality. Quantitative evaluations in Tab. 3 reveal that PSHuman outperforms existing methods across multiple metrics, achieving the highest PSNR ($20.8548$), SSIM ($0.8636$) as well as the lowest LPIPS ($0.0764$), which correlates more closely with visual perception. Qualitatively, as illustrated in Fig. 7, PSHuman produces highly consistent appearances on novel viewpoints, including natural and realistic reconstruction for posterior regions. In contrast, existing methods exhibit various limitations such as blurred colors and inconsistent artifacts in unseen views.
Robustness to SMPL-X estimation. We assess the robustness of template-based approaches to SMPL-X estimation errors in Tab. 3. Following SIFU, we introduce random noise with a variance of $0.05$ to both the pose and shape parameters of the ground-truth SMPL-X model. The results demonstrate the robust reconstruction capabilities of our approach. Furthermore, the efficacy of our method in real-world scenarios is evidenced by the additional results presented in Fig. 13 of  A.4.
@@ 4.2 Ablation Study
Effectiveness of SMPL-X condition. In Fig. 3, we show the geometry reconstructed by the models trained without SMPL-X condition and with SMPL-X condition. In Fig. 3(a), it is observed that the naive diffusion model struggles to ’imagen’ the pose of a bending human image. Conversely, the SMPL-X provides a strong pose prior to guide the model to generate reasonable side views, leading to better reconstruction. In Fig. 3(b), the diffusion model fails to generate consistent multiple views due to self-occlusion, resulting in artifacts near the art regions. The SMPL-X guidance effectively enhances consistency, facilitating the complete human body.
     Input           w/o CSD                 w/ CSD
@@
Effectiveness of cross-scale diffusion (CSD). In Fig. 8, we experiment with the removal of the locally enhanced model, which means only usage of the global diffusion branch. The resulting appearance and geometry, as can be observed, are obviously distorted (e.g. the mouth region) or blurry and fail to accurately recover the consistent geometry details with reference input image. However, using the local enhanced diffusion model, our method manages to overcome these limitations. It achieves more precise and intricate details, contributing to a significant enhancement for the appearance and geometry of 3D humans.
Figure 9: Ablation of our reconstruction module.
@@
Effectiveness of mesh carving module. We assess the efficacy of our reconstruction module by substituting the remeshing step with alternative methods, specifically NeuS and BiNI. As illustrated in Fig. 9, the resulting geometries exhibit notable deficiencies or failures to capture fine geometric details. Note that we employ the normal maps, generated by our diffusion model, across all methods to mitigate potential errors arising from normal prediction discrepancies. Moreover, in the absence of SMPL-X optimization, the reconstructed mesh displays subtle artifacts due to misalignment between the initial SMPL-X and the multiple views. Our reconstruction module, which incorporates remeshing with SMPL-X refinement, effectively addresses these issues. For a comprehensive evaluation, we direct the reader to Sec. A.4.
5 Limitations and Conclusion
In this work, we present PSHuman, a single-view human reconstruction framework that significantly enhances the quality of both geometry and appearance. By introducing a body-face cross-scale diffusion model, we improve the capability of modeling high-fidelity 3D human faces. Additionally, we use SMPL-X as guidance for robust multi-view generation. Finally, we devise the multi-view guided explicit human carving module to preserve as many details from generated images as possible. We demonstrate that PSHuman can generate 3D humans with intricate geometric details and realistic appearances, outperforming existing methods.
Limitations. We share a common problem with previous template-based works: the pose estimation error has a cascading effect on subsequent view generation and reconstruction. It is promising to mitigate it by unifying existing multi-view datasets and improving the generation robustness without body template conditions.
Appendix A Appendix
A.1 Discussions about face-body cross-scale diffusion
Difficulty in implementing dependent forward process. In the dependent forward process $q(x^{B}_{t}|x^{B}_{t-1},x^{F}_{t-1})$, we know that the face region of $x^{B}$ corresponds to $x^{F}$. Since we have defined $p(x^{F}_{t}|x_{t-1}^{F})$ by adding noises to $x^{F}_{t-1}$, it is natural to get $x^{B}_{t}$ by replacing the pixel values in the face region of $x^{B}_{t}$ with $x^{F}_{t}$ and just adding noises to the remaining image regions of $x^{B}_{t-1}$. However, since we adopt a latent diffusion model (Stable Diffusion) Rombach et al. (2022a) here, the pixels of tensors in the latent spaces are not independent of each other so the replacing operation is not valid here. This brings difficulty in separating the face regions in the latent space to explicitly implement the dependent forward process for adding noises.
A.2 Implementation Details
Preprocessing. Our training datasets include scans from THuman2.1 and CustomHumans. For each human model, and the corresponding SMPL-X model, we render 8 color and normal images with alpha channel around the yaw axis, with a $45^{\circ}$ interval and a resolution of $768\times 768$. Due to the random face-forward direction, we employ insightface Deng et al. (2018) for face detection, utilizing only viewpoints containing clear facial characteristics for training. As mentioned in the main paper, PSHuman generates 6 color and normal images from front, front-right, right, back, left, and front-left views. To guarantee the generation alignment, we horizontally flip the left and back views during training.
Diffusion block. As shown in Fig. 5, our diffusion block consists of two branches, in which the local diffusion inherits from stable diffusion, including self attention, cross attention and feed-forward layers, while the global attention contains an additional multi-view attention layer introduced in Era3D. Global attention is conditioned on the local branch via the alignment of hidden layers.
Training and evaluation details. PSHuman builds upon the open-source pre-trained text-to-image generation model, SD2.1-unclip Rombach et al. (2022b). Our training is conducted on a cluster of 16 NVIDIA H800 GPUs, with a batch size of 64 for a total of 30,000 iterations. We adopt an adaptive learning rate schedule, initializing the learning rate at 1e-4 and decreasing it to 5e-5 after 2,000 steps. The entire training process spans approximately 3 days. To enable class-free guidance (CFG) Ho & Salimans (2022) during inference, we randomly omit the clip condition at a rate of 0.05 during training. During inference, we employ PyMAF-X Zhang et al. (2023) for hand pose estimation and PIXIE Feng et al. (2021) for body pose prediction for robustness. For the reconstruction module, we set the number of steps for SMPL-X alignment, geometry optimization, and texture fusion to 700, 100, and 100, respectively, with corresponding learning rates of 0.3, 0.001, and 0.0005. Regarding appearance evaluation, we render color images from four viewpoints at azimuths of $0^{\circ},90^{\circ},180^{\circ},270^{\circ}$ relative to the input view.
Table 4: Inference time of the reconstruction module.
Pipeline Pre-processing Diffusion Geo. Recon. Appearance Fusion
Time / s 7.2 17.6 23.3 6.0
A.3 User study
Given the limitations of quantitative metrics in assessing the realism and consistency of side and back views reconstructed from single-view input, we conducted a comprehensive user study to evaluate the geometry and appearance quality of five SOTA methods.
We collect 20 in-the-wild samples and 20 cases from SHHQ fashion dataset for evaluation. Following HumanNorm Huang et al. (2024a), we invite 20 volunteers to evaluate the color and normal video rendered from the reconstructed 3D humans. Participants were instructed to score each model on a 5-point scale (1 being the worst and 5 being the best) across four key dimensions:
  • To what extent does the human model exhibit the best geometry quality?
Method PIFuHD PaMIR ECON GTA SiTH Ours
Geometry Quality 1.55 1.96 3.72 2.11 2.72 4.71
Appearance Quality - 1.42 - 2.65 2.82 4.59
Geometry Consistency 1.69 1.76 2.48 2.33 2.79 4.61
Appearance Consistency - 1.77 - 2.16 2.73 4.68
A.4 More experiments
Comparison with optimization-based methods. To assess the efficacy of our approach relative to optimization-based methods, we conducted a comparative analysis of PSHuman against several SDS-based techniques, Magic123, Dreamgaussian, Chupa, and TeCH. Following SiTH, we adopt the pose and text prompt generated by (Li et al., 2022) as condition inputs due to the lack of direct image input support in Chupa. As illustrated in Fig. 10, Magic123 and Dreamgaussian exhibit significant limitations, primarily manifesting as incomplete human body reconstructions and implausible free-view textures. The reliance on text descriptions for conditioning proves insufficient for fine-grained control, resulting in geometries that deviate substantially from the reference inputs. TeCH, a method specifically designed for human reconstruction from a single image, while capable of producing complete human shapes, struggles with severe noise in geometric details and over-saturated textures. These artifacts are characteristic challenges inherent to SDS-based methodologies. In contrast, PSHuman demonstrates superior performance by directly fusing multi-view 2D images in 3D space, enabling the preservation of geometry details at the pixie level while circumventing unrealistic texture. Note that TeCH requires $\sim$6 hours for optimization, PSHuman generates high-quality textured meshes within merely 1 minute.
      Input       view=2           view=4            view=6 (Ours)
@@ A.4 More experiments
Comprehensive quantitative ablation. In addition to the qualitative ablation in Fig. 3 and Fig. 8, we further conducted comprehensive ablation studies on a subset of 20 samples from the ”CAPE-NFP” dataset. Tab. 6 quantitatively illustrates the impact on Chamfer Distance performance when individual components are removed or replaced. It is observed that the SMPL-X condition contributes significantly to reconstruction accuracy. While CSD yields a modest reduction in geometric error, it substantially improves visualization quality and identity fidelity, as evidenced in Fig .8. Furthermore, our reconstruction method, which employs SMPLX-guided differentiable remeshing, demonstrates superior reconstruction performance compared to the BiNI and inpainting pipeline utilized in ECON. The overall results showcase the efficacy of each component in achieving high-quality 3D human reconstruction.
@@
Table 6: The ablation study of core designs.

### v2

@@ 3.2 SMPLX-initialized Explicit Human Carving
Following the generation of multiview color and normal images, we elaborate on our SMPLX-initialized human carving module (Fig. 3) to obtain the textured 3D mesh.
@@
Numerous methodologies have been developed to leverage normal cues for human reconstruction. However, a significant proportion of them employ implicit functions (e.g. MLP) to map the normal feature as implicit surfaces. This process, while effective in certain scenarios, often results in a lack of fine geometric details. Even with BiNI used in ECON, the overall geometry still exhibits a notable degradation. Taking advantage of the multiview consistent normal maps, we opt to fuse it directly with the explicit triangle mesh. Our reconstruction module consists of three main stages: SMPL-X initialization, differentiable remeshing, and appearance fusion.
@@
SMPL-X initialization. The process commences with human mesh initialization, utilizing the aforementioned SMPL-X estimation, which provides a strong body prior, effectively mitigating unnecessary face pruning and densification during subsequent geometry optimization. However, it is noteworthy that the generated multiple views may exhibit slight misalignment with the SMPL model due to normalization and recentering procedures. Drawing inspiration from ICON, we optimize SMPL-X’s translation $\tilde{t}$, shape $\tilde{\beta}$, and pose $\tilde{\theta}$, parameters to minimize:
$\tilde{t},\tilde{\beta},\tilde{\theta}=\operatorname*{arg\,min}_{t,\beta,\theta}\sum_{i=1}^{N}w_{i}(\|N_{i}-\hat{N}_{i}\|_{2}+\|S_{i}-\hat{S}_{i}\|_{2}),$ (7)
where $w_{i}$ denotes the confidence of $i$-th view, $\hat{N}_{i}$ and $\hat{S}_{i}$ represent the SMPL-X normal and silhouette renderings from predefined views.
@@
Remeshing with differentiable rasterization. Given the initial human prior, we utilize differentiable rasterization to carve the details based on observational normal maps. While a common approach involves adding per-vertex displacement to the coarse canonical mesh, this method encounters difficulties when modeling complex details, such as loose clothing. To address this limitation, we directly optimize the SMPL topology, encompassing both vertex positions $V$ and face edges $F$. The optimization procedure iteratively applies vertex displacement and remeshing to the triangle mesh, utilizing the optimizer proposed in [30]. The optimization objective can be written as
$\tilde{V},\tilde{F}=\operatorname*{arg\,min}_{V,F}\sum_{i=1}^{N}w_{i}(\|N_{i}-\hat{N}_{i}\|_{2}+\|S_{i}-\hat{S}_{i}\|_{2})+\lambda\sum_{j}(n_{j}-n_{j}^{\text{neig}}),$ (8)
@@
where $w_{i}$ denotes the confidence of $i$-th view, $\hat{N}_{i}$ and $\hat{S}_{i}$ represent the normal and silhouette renderings from predefined views, $n_{j}$ and $n_{j}^{\text{neig}}$ denote the vertex normal and the average normal of neighboring vertices. The regularization weight $\lambda$ is set to $0.02$. We execute $700$ optimization steps to achieve optimal performance. Following the mesh optimization, we employ Poisson reconstruction [17] to complete minor invisible areas, such as the chin. Additionally, following [43], we offer the option to replace the hands with the estimated SMPL-X results to enhance visual quality.
@@
Appearance fusion. Upon obtaining the 3D geometry, our objective is to derive the high-fidelity texture matching the reference image. Despite the availability of multiview images, direct projection onto the mesh results in conspicuous artifacts, arising from the cross-view inconsistency and inaccurate foreground segmentation. To overcome this, we perform texture fusion utilizing the aforementioned differentiable rendering. Specifically, we optimize the per-vertex color $VC$ by minimizing
$VC=\arg\min_{vc}\sum_{i=1}^{N}w_{i}\|C_{i}-\hat{C}_{i}\|_{2},$ (9)
where $C_{i}$ and $\hat{C}_{i}$ represent the rendered and generated color images, respectively. In the majority of cases, this color fusion pipeline suffices to generate high-quality appearances. However, certain areas may remain unobserved from the predefined six viewpoints. Thus, we finally compute a visibility mask and perform topology-aware interpolation based on KDTree, ensuring comprehensive texture coverage.
@@ 4 Experiments
   Input     PaMIR      PIFuHD         GTA         SiFU        HiLo        ECON          Ours
Figure 4: Geometry comparison of PSHuman with Implicit and Explicit methods for 3D human inference from in-the-wild images. Existing methods often struggle with complex poses and loose clothing, leading to issues such as absent body parts, disrupted clothing, and a lack of fine details. In contrast, PSHuman provides a complete shape, detailed facial features, and natural-looking clothing folds. Following [43], we substitute the hands with SMPL-X models to enhance visual quality.
Training and evaluation details. PSHuman builds upon the open-source pre-trained text-to-image generation model, SD2.1-unclip [34]. Our training is conducted on a cluster of 16 NVIDIA H800 GPUs, with a batch size of 64 for a total of $30,000$ iterations. We adopt an adaptive learning rate schedule, initializing the learning rate at 1e-4 and decreasing it to 5e-5 after $2,000$ steps. The entire training process spans approximately three days. Regarding the reconstruction module, we perform SMPL-X alignment, geometry optimization, and texture fusion for $700$, $100$ and $100$ steps, respectively, with corresponding learning rates of $0.3$, $0.001$, and $0.0005$. For appearance evaluation [51], we render color images from four viewpoints at azimuths of {$0^{\circ},90^{\circ},180^{\circ},270^{\circ}$} relative to the input view.
Dataset. We conduct experiments on widely used 3D human datasets, including THuman2.1 [47], CustomHumans [12] and CAPE [25]. Specifically, our training dataset comprises $2,385$ scans from THuman2.1 and $647$ scans from CustomHumans. These datasets are selected due to their provision of SMPL-X parameters. For quantitative evaluation, we utilize the remaining $60$ scans ($0447$-$0486$, $0492$-$0511$) from THuman2.1 and $150$ scans from CAPE. Following ICON’s partitioning criteria, we subdivide CAPE into ”CAPE-FP” (50 samples) and ”CAPE-NFP” (100 samples) to assess generalization in real-world scenarios.
Metric. To assess reconstruction capability, we employ three primary metrics: 1-directional point-to-surface (P2S), $L_{1}$ Chamfer Distance (CD), and Normal Consistency (NC). For appearance evaluation, we utilize peak signal-to-noise ratio (PSNR[41], structural similarity index (SSIM[48], and learned perceptual image patch similarity (LPIPS[49].
Table 1: Quantitative comparison of geometry quality. For the setting of ‘w/o SMPL-X body prior’, we utilize PIXIE to estimate SMPL parameters for other baseline methods while omitting SMPL estimation for our approach. Specifically, we retrain the diffusion model by removing the SMPL-X conditioning and initialize human mesh with a unit sphere during mesh carving. For ‘w/ SMPL-X body prior’, ground-truth SMPL-X models are used to avoid the impact of pose estimation errors on the evaluation. The units for Chamfer and P2S are in cm. The top two results are colored as first second.
@@
Method Venue Cham. Dist $\downarrow$ P2S $\downarrow$ NC $\uparrow$ Cham. Dist $\downarrow$ P2S $\downarrow$ NC $\uparrow$ Cham. Dist $\downarrow$ P2S $\downarrow$ NC $\uparrow$
@@ 4.1 Comparisons
Baselines. We conducted a comprehensive comparison of our method against state-of-the-art single-view human reconstruction approaches, including PIFu [35], PIFuHD [36], PaMIR [54], ICON [42], ECON [43], GTA [50], SiFU [51], HiLo [45], and SiTH [13]. For SMPL-based methods, we utilize PIXIE [46] for estimation. We also report the results with ground-truth SMPL-X to isolate the impact of pose estimation errors.
Comparison of geometry quality. Leveraging consistent multiview images, our method exhibits superior geometric quality compared to existing approaches, particularly in scenarios without SMPL-X body prior (Tab. 1). Unlike other template-based methods, which are susceptible to SMPL-X prediction errors, our method supports template-free training, thereby offering enhanced generalization capability. When incorporating the body prior, our method consistently outperforms previous works, demonstrating unprecedented accuracy on complex posed humans. The qualitative comparison in Fig. 4 also showcases the superiority of PSHuman, featuring with complete shape, detailed face and natural-looking clothing folds.
Table 2: Quantitative comparisons of face reconstruction.
Method Cham. Dist $\downarrow$ P2S $\downarrow$ NC $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
ECON 0.624 0.570 0.837 - - -
SIFU 0.535 0.527 0.853 18.86 0.790 0.093
SITH 0.610 0.563 0.858 17.93 0.827 0.110
w/o local 0.524 0.503 0.867 19.67 0.832 0.093
w/o noise blender 0.447 0.422 0.904 20.85 0.877 0.075
Ours 0.423 0.397 0.924 20.97 0.896 0.071
@@
Comparison of appearance quality. Quantitative evaluations in Tab. 4 reveal that PSHuman outperforms existing methods across multiple metrics, achieving the highest PSNR, SSIM as well as the lowest LPIPS. Qualitatively, as illustrated in Fig. 5, PSHuman produces highly consistent appearances on novel viewpoints, including natural and realistic reconstruction for posterior regions. In contrast, existing methods exhibit various limitations such as blurred colors and inconsistent artifacts in unseen views.
Comparisons of face quality. To highlight the effectiveness of our introduced cross-scale diffusion for face reconstruction, we use the head vertices of SMPL-X to crop the reconstructed head following ECON. Specifically, we first construct a KD-tree based on SMPL-X to query the generated mesh, subsequently filtering out the vertices adjacent to the head of SMPL-X. Tab. 2 presents the quantitative comparisons with SOTA methods.
@@ 4.2 Ablation Study
Effectiveness of SMPL-X condition. In Fig. 1, we show the geometry reconstructed by the models trained without SMPL-X condition and with SMPL-X condition. In Fig. 1(a), it is observed that the naive diffusion model struggles to ‘imagine’ the pose of a bending human image. Conversely, the SMPL-X provides a strong pose prior to guide the model to generate reasonable side views, leading to better reconstruction. In Fig. 1(b), the diffusion model fails to generate consistent multiple views due to self-occlusion, resulting in artifacts near the arm regions. The SMPL-X guidance effectively enhances consistency, facilitating the complete human body.
      Input      w/o local   w/o noise blendering   Ours
@@
Effectiveness of cross-scale diffusion (CSD). In Tab. 2, we provide the results by removing the local face branch (w/o local) and noise blending (w/o noise blending), respectively. Our method, incorporating both components, achieves the highest performance, as shown in Fig. 6. Notably, the setting without noise blending also generates the local face image. However, the reconstructions exhibit minor artifacts or over-smoothness. We attribute it to the inconsistency among global and local images. In contrast, the noise blending allows the information exchange, thereby enhancing overall consistency.
    Input      NeuS     BiNI     SMPLX-D  Naive Remesh  Ours
Figure 7: Ablation of our human carving module.
Figure 8: Visualization of mesh carving of a posed human image.
@@
Effectiveness of mesh carving module. We assess the efficacy of our reconstruction module by substituting the remeshing step with alternative methods, specifically NeuS and BiNI. As illustrated in Fig. 7, the resulting geometries exhibit notable deficiencies or failures to capture fine geometric details. Note that we employ the normal maps, generated by our diffusion model, across all methods to mitigate potential errors arising from normal prediction discrepancies. Moreover, “naive remeshing” refers to remeshing with SMPL-X initialization but without multiview-guided SMPL-X alignment, resulting in subtle artifacts caused by misalignment between the initial SMPL-X mesh and the multi-view observations. Our reconstruction module effectively addresses these issues. Finally, we show an example across remeshing process for better understanding in Fig. 8.
Robustness to SMPL-X estimation. We assess the robustness of template-based approaches to SMPL-X estimation errors in Tab. 4. Following SIFU, we introduce random noise with a variance of $0.05$ to both the pose and shape parameters of the ground-truth SMPL-X model. The results demonstrate the robust reconstruction capabilities of our approach. Furthermore, the efficacy of our method in real-world scenarios is evidenced by the additional results presented in supplementary materials.
@@
Comprehensive quantitative ablation. We further conducted comprehensive ablation studies on a subset of $20$ samples from the “CAPE-NFP” dataset. Tab. 5 quantitatively illustrates the impact on Chamfer Distance when individual components are removed or replaced. It is observed that the SMPL-X condition contributes significantly to reconstruction accuracy. While CSD yields a modest reduction in geometric error, it substantially improves visualization quality and identity fidelity, as evidenced in Fig. 6. Furthermore, our reconstruction method, which employs SMPLX-guided differentiable remeshing, demonstrates superior reconstruction performance compared to the BiNI and inpainting pipeline utilized in ECON.
@@
Table 5: Comprehensive ablation study of core designs w.r.t full body reconstruction performance.
@@
5 Conclusion
A. Inaccurate pose B. Stitching artifacts C. Wrong generation    D. Loose hair
Figure 9: Failure cases of PSHuman.
Limitations. Although PSHuman achieves high-quality single-view human reconstruction, it shares certain limitations with previous template-based approaches as shown in Fig. 9. First, pose estimation errors (A, B) have a cascading effect on subsequent generation and reconstruction, impacting overall accuracy. In addition, wrong novel-view generation (C) may result in unreasonable geometry. Finally, diffusion models struggle to generate consistent subtle details, such as loose hair and hands (D), which results in suboptimal reconstruction.
Conclusion. We present PSHuman, a novel framework that significantly improves geometric and appearance quality in single-image human reconstruction. We investigate direct multiview human generation conditioned on SMPL-X, enabling explicit and robust human reconstruction. Our body-face cross-scale diffusion model enhances the modeling of high-fidelity 3D human faces, while our multiview-guided explicit carving module ensures intricate details from generated images. Experiments demonstrate that PSHuman’s superiority against existing methods.
Acknowledgement. The research was supported by Theme-based Research Scheme (T45-205/21-N) from Hong Kong RGC, Generative AI Research, Development Centre from InnoHK, NSFC (No.62171255) and Guoqiang Institute of Tsinghua University (No.2021GQG0001).
Supplementary Material
6 Discussions about face-body cross-scale diffusion
Difficulty in implementing dependent forward process. In the dependent forward process $q(x^{B}_{t}|x^{B}_{t-1},x^{F}_{t-1})$, we know that the face region of $x^{B}$ corresponds to $x^{F}$. Since we have defined $p(x^{F}_{t}|x_{t-1}^{F})$ by adding noises to $x^{F}_{t-1}$, it is natural to get $x^{B}_{t}$ by replacing the pixel values in the face region of $x^{B}_{t}$ with $x^{F}_{t}$ and just adding noises to the remaining image regions of $x^{B}_{t-1}$. However, since we adopt a latent diffusion model (Stable Diffusion) Rombach et al. [33] here, the pixels of tensors in the latent spaces are not independent of each other so the replacing operation is not valid here. This brings difficulty in separating the face regions in the latent space to explicitly implement the dependent forward process for adding noises.
7 Implementation Details
Preprocessing. Our training datasets include scans from THuman2.1 and CustomHumans. For each human model, and the corresponding SMPL-X model, we render 8 color and normal images with alpha channel around the yaw axis, with a $45^{\circ}$ interval and a resolution of $768\times 768$. Due to the random face-forward direction, we employ insightface Deng et al. [5] for face detection, utilizing only viewpoints containing clear facial characteristics for training.